Estimate the total hydrostatic force ? on the Hoover Dam.

Hint: Look up the dimensions of the Hoover Dam, the elevation above sea level of its crest, and the current elevation of the water level in Lake Mead.

Hoover Dam dimensions: 890,000 m^2

Elevation Crest: 376 m
Answer #1

Hydrostatic force on dam,

F = P*A

F = (\rho*g*d/2 ) * A

F = (1000*9.8*376 / 2) * 890000

F = 1.64*1012 N